OpenStudy (anonymous):

The ratio of the length of an airplane wing to its width is 8 to 1. if the length of a wing is 32.7 meters how wide must it be

OpenStudy (anonymous):

\[\large 8~to~1~=~\frac{8}{1}=8\]\[\large 32.7~to~w~=~\frac{32.7}{w}\]\[\large 8~=~\frac{32.7}{w} \Rightarrow 8w=32.7 \Rightarrow w=\frac{32.7}{8}\]